OpenStudy (anonymous):

Will give Medal and Fan x + 15 = 8 x = ?

OpenStudy (ahsome):

Equation is: \[x+15=8\] Move the 15 to the RHS \[x=8-15\]\[x=-7\] Hope that helps :)

OpenStudy (anonymous):

thanx

OpenStudy (ahsome):

No problem :D

OpenStudy (anonymous):

now would you work 12 - x = 7 the same way

OpenStudy (jessiegonzales):

12-7=x x=5

OpenStudy (anonymous):

now how would you work that